OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3.

OpenStudy (anonymous):

@easyaspi314

OpenStudy (anonymous):

Slope intercept equation is y = mx+b

OpenStudy (anonymous):

so 1 = -2x + 3 i think :P

OpenStudy (shamil98):

point-slope form is y - y1 = m(x- x1) y - 1 = 3(x+2)

OpenStudy (anonymous):

lmaoo, im not that gud at math @shamil98

OpenStudy (shamil98):

its just plugging into a formula lol.

OpenStudy (anonymous):

lmaoo, i didnt know the formula, i had the wrong one :P

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

can someone help me understand how to do thi problem? I have to do a test an have no idea how to do this!

OpenStudy (anonymous):

get the formula and plug in wat was given to u into the formula :)

OpenStudy (shamil98):

The point-slope formula is \[y - y_1 = m(x-x_1)\] just plug in the x y coordinates (x,y) and the slope. :P

OpenStudy (anonymous):

exactly, lmaoo

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

good job @shamil98 :) Good luck @kcurri

OpenStudy (anonymous):

thanks